OpenStudy (anonymous):

Derivative of tanx cos^2 x

OpenStudy (anonymous):

@IMStuck @ganeshie8

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

(A) 1+ 2sin^2 x (B)1 - 2sin^2 x (C) 1 (D) 2sin^2 x

OpenStudy (anonymous):

GOT IT !!!!!!

OpenStudy (tkhunny):

Did you make your life simple or do it the hard way? \(\tan(x)\cdot\cos^{2}(x) = \sin(x)\cos(x) = ½\sin(2x)\) Note: Away from \(\cos(x) = 0\), anyway.

OpenStudy (anonymous):

Made my life harder i opened it using product rule

OpenStudy (anonymous):

everything cancels out and you get (B)

OpenStudy (tkhunny):

Well, there. Now you have two additional expressions to try to see if you get the same thing. Don't forget your trigonometry! It is hoped that you find B is also \(\cos(2x)\). :-)

OpenStudy (anonymous):